OpenStudy (anonymous):

(Basic stats): If n=9, p=.333... (see full question)

OpenStudy (anonymous):

In order to test people's discrimination abilities, or their ability to tell things apart, a simple experiment is planned consisting of 9 trials. In this particular experiment, participants are asked whether they can differentiate between three major cola brands. On each trial, participants must discriminate between the three alternatives, that is, they are asked to guess which brand of cola they are drinking. Suppose participants cannot tell the brands apart. 1. Calculate the exact probability of participants being wrong on 2 or more trials. 2. Calculate the exact probability of participants being wrong on ALL trials.

OpenStudy (anonymous):

all trials is easiest

OpenStudy (anonymous):

probability you are wrong is \(\frac{2}{3}\) i guess, making the probability you are wrong 9 times in a row is \[\left(\frac{2}{3}\right)^9\]

OpenStudy (anonymous):

i hate word problems, don't you?

OpenStudy (anonymous):

Yes, I do!! Thank you!